OpenStudy (anonymous):

Write the equation for this situation: the cost, c, in dollars of a car rental is equal to 10 dollars plus 1/2 the miles driven, m

OpenStudy (anonymous):

c=10+m/2

OpenStudy (anonymous):

The cost C is equal to 10 dollars + half of m Mathematically, c=10+m/2 Taking 2 to the left hand side we have 2C=10+M Take different values of C and M. Assume C to be X and M to be Y. You will get different values of c for different values of m. Plot these co-ordinate points and join them to make a graph. If the number of miles driven is 12, the cost would be: 2C=10+12 2C=22 C=22/2 C=$11

OpenStudy (anonymous):

thanx